# Computing the number of realisations of a rigid graph

### v1

@@ 1.2 Statement of result
Given Theorem 1, we can calculate the realisation number for any rigid graph. Indeed, let $G$ be any rigid graph. If $G$ is minimally rigid, then by the existing algorithm of [CGG+18], its realisation number can be computed. If $G$ is globally rigid, its realisation number is $1$. If neither of these occur, we apply Theorem 1; we can write $c_{2}(G)$ in terms of realisation numbers of graphs on strictly fewer vertices. The algorithm then proceeds recursively; if any of the smaller graphs are minimally rigid or globally rigid, we can compute their realisation number. If not, we apply Theorem 1 again. This process clearly concludes at some point, since we have a finite number of vertices in $G$ and the number of vertices in the graphs considered strictly decreases at each stage.
It is natural to ask how often Theorem 1 is needed when computing realisation numbers. Table 1 shows how many graphs on few vertices are minimally/globally/redundantly rigid. What is particularly interesting is the last column of Table 1, which shows the number of rigid graphs which are neither minimally rigid nor redundantly rigid and which do not have a 2-cut, i. e., those graphs that require Theorem 1(i) (the more computationally involved of the two cases) to compute their realisation number.
@@ 1.3 Examples
Figure 1: (Left): The rigid graph $G$ given in Example 2. (Middle): The maximally rigid subgraphs of $G^{\prime}=G-\{1,2\}$. (Right): The minimally rigid subgraph $H$ of $G$.
@@
Let us consider the graph $G$ from Figure 1 (left). This graph is clearly not minimally rigid since it has one too many edges. It is not redundantly rigid either and hence not globally rigid, since the deletion of the edge $\{1,2\}$ would result in a non-rigid graph $G^{\prime}$.
We split the graph $G^{\prime}$ into its maximal rigid components $G_{1},\ldots,G_{6}$ as shown in Figure 1 (middle). Let $G_{1}$ be the graph with vertices $4,5,6,7$. It is the only subgraph that is not minimally rigid. We compute $c_{2}(G_{1})=1$ since it is globally rigid. Each other subgraph $G_{i}$ is a single edge, and hence the respective $H_{i}$ are also single edges. This in turn implies $c_{2}(G_{i})=c_{2}(H_{i})=1$ for each $i\neq 1$. It is not so hard to convince oneself that the graph $H_{1}$ obtained from $G_{1}$ by deleting the edge $\{5,7\}$ is minimally rigid with $c_{2}(H_{1})=2$. The algorithm given in [CGG+18] applied to the graph $H=H_{1}\cup\ldots\cup H_{6}\cup\{1,2\}$ gives that $c_{2}(H)=24$. By applying Theorem 1(i) we get $c_{2}(G)=c_{2}(H)\prod_{i=1}^{m}\frac{c_{2}(G_{i})}{c_{2}(H_{i})}=24\cdot\frac{1}{2}=12$.
@@
Now consider the graph $\widetilde{G}$ from Figure 2 (left). Again, this graph is neither minimally rigid nor is it redundantly rigid. The algorithm described in [CGG+18] gives that the graph $\widetilde{H}=\widetilde{G}-\{5,7\}$ has a realisation number of 672. By combining Theorem 1(i) ($e=\{10,11\}$) with the observation that the subgraph generated on the first 7 vertices is the graph $G$ from Example 2, we have that $c_{2}(\widetilde{G})=c_{2}(\widetilde{H})\cdot\frac{c_{2}(G)}{c_{2}(H)}=672\cdot\frac{12}{24}=336$.
@@
Figure 2: (Left): The rigid graph $\widetilde{G}$ given in Example 3. (Middle): The maximally rigid subgraphs of $\widetilde{G}-\{10,11\}$. (Right): The minimally rigid subgraph $\widetilde{H}$ of $\widetilde{G}$.
@@ 1.4 Structure of paper and notation
This paper is structured as follows. We cover all the required background material for rigidity theory in Section 2. In Section 3, we provide a proof of Theorem 1. This algorithm is then adapted in Section 4 to compute spherical realisation numbers and rank-3 PSD matrix completions. We conclude the paper in Section 5 with a range of computational results.
@@ 2.2 Algebraic geometry and rigidity
We now wish to view our realisation numbers as degrees of dominant maps of varieties. We note that different authors use various definitions of a ‘realisation variety’, using techniques such as pinning vertices (e.g., [JO19, DG24]). In this paper we do not wish to use pinning, as it makes the upcoming decomposition map (Section 3) difficult to define. Instead, we want to exploit the following useful property of the variety $R_{V}$: by projecting away coordinates we can reach $R_{V^{\prime}}$ for any $V^{\prime}\subseteq V$.
@@
If $G$ is rigid, then the number of connected components of $f_{G}^{-1}(f_{G}(\rho))$ for generic $\rho$ is $2c_{2}(G)$; see, for example, [LMSW25, Lemma 5]. Similarly, if $G$ is rigid then $\deg(p_{G})=c_{2}(G)$. A proof of this fact can be found in Appendix A.
3 Proof of Theorem 1
As stated previously, Theorem 1(ii) was originally proven by Jackson and Owen [JO19, Theorem 6.6]. We now proceed with Theorem 1(i). The main idea of the proof is to factorise the graph map $p_{G}$ in a suitable way, using the following map.
@@ 3 Proof of Theorem 1
The proof of Lemma 7 extends to higher dimensions, however the decomposition of $H$ into minimally rigid subgraphs sharing at most one vertex is no longer guaranteed. This is important for the dominance of the decomposition map $d_{S}$; if, in three dimensions, two of the subgraphs, say $H_{1}$ and $H_{2}$ of $G$ share a pair of vertices $v_{1},v_{2}$, then after applying $d_{S}$ to any realisation $\rho$ of $G$, we must have that, for $d_{S}(\rho)=(\rho_{1},\rho_{2},\ldots,\rho_{t})$, the equation $d(\rho_{1}(v_{1}),\rho_{1}(v_{2}))=d(\rho_{2}(v_{1}),\rho_{2}(v_{2}))$ is always satisfied. This shows that $d_{S}$ cannot be a dominant map in such a situation. The dominance of $d_{S}$, however, is crucial for the proof of Theorem 1. Hence, extending our results to higher dimensions is a subject for further research.
@@
The inclusion $I_{G}\subseteq I_{G_{1}}\times\ldots\times I_{G_{m}}\times\mathbb{C}$ is clear. It now suffices to show that both sets have the same dimension: since both $I_{G}$ and $I_{G_{1}}\times\ldots\times I_{G_{m}}\times\mathbb{C}$ are irreducible and closed in the Zariski topology, this immediately implies the sets are equal. Since $G$ is rigid, we have $\dim I_{G}=r(G)=2|V(G)|-3$. By Lemma 6, we then have
@@
We are now ready to prove the first case of Theorem 1.
@@
As in the proof of Lemma 7, the graph map $p_{G}:R_{V(G)}\to I_{G}$ factors into the product of graph maps of subgraphs
@@
and the decomposition map $d_{S}$; this can be seen in the following commutative diagram, where we note that by Lemma 9 we have $I_{G}=I_{G_{1}}\times\cdots\times I_{G_{m}}\times I_{e}$:
@@
The map $d_{S}$ is also the decomposition map for the decomposition $\{H_{1},\dots,H_{m},e\}$ of $H$, since $R_{V(G)}=R_{V(H)}$ and $R_{V(G_{i})}=R_{V(H_{i})}$ for all $i$. By Lemma 5, $H$ is minimally rigid, and by Lemma 7, $d_{S}$ is a generically finite and dominant map with degree $\frac{c_{2}(H)}{c_{2}(H_{1})\cdots c_{2}(H_{m})}$. For the subgraph $e$, note that $R_{e}=I_{e}=\mathbb{C}$, and $p_{e}$ is the identity map. We then have
@@
Proof of Theorem 1.
Let $G$ be a graph that is rigid but not globally rigid. By [Hen92], $G$ is either not redundantly rigid, or $G$ is not 3-connected. If $G$ is either not redundantly rigid, then there exists an edge $e\in E(G)$ such that GeG-e is not rigid. Case (i) now follows by Lemma 10. If $G$ is not 3-connected, we see that Case (ii) holds by [JO19, Theorem 6.6]. ∎
@@ 4.1 Computing spherical realisation numbers
Our proof of Theorem 1(i) also extends to counting realisations on the (complex) sphere $\mathbb{S}^{2}$. Indeed, a graph is minimally rigid in $\mathbb{C}^{2}$ if and only if is also minimally rigid on the sphere (see for instance [SW07]). Furthermore, a graph is globally rigid in $\mathbb{C}^{2}$ if and only if it is globally rigid in $\mathbb{S}^{2}$; the real variant of this statement follows from a result of Connelly and Whiteley [CW10, Theorem 12], which can then be adapted to the complex setting using a combination of results from Gortler and Thurston [GT14, Theorem 1] and the first and second authors [DG24, Theorem 1.2]. Therefore, Jackson and Jordán’s characterisation of globally rigid graphs also holds in the sphere, and we have the same division into two cases.
@@
The graph from Example 2 has $c^{\circ}_{2}(G)=16$ since $c^{\circ}_{2}(H)=32$. This is the smallest rigid but not minimally rigid graph where $c_{2}(G)\neq c^{\circ}_{2}(G)$.
@@ 5.1 Computational results for realisation numbers
We focus now on rigid graphs with 2|V|3+k2|V|-3+k edges and a high number of realisations. For $k=0$, i.e. the minimally rigid case, this was done before in [GKT20, Gra25]. Let $\mathbf{M}_{2}^{k}(n)$ be the maximal $c_{2}(G)$ over all rigid graphs $G=(V,E)$ with |E|=2|V|3+k|E|=2|V|-3+k. We show in Table 2 some values of $\mathbf{M}_{2}^{k}(n)$, which can be found by computing the realisation numbers of all rigid graphs with the respective number of vertices and edges. Note that for $n\leq 10$ we have computed realisation numbers of all rigid graphs which satisfy the desired edge count. They can be found at [DGS+25]. For $n\geq 11$ we only computed realisations numbers for all rigid graphs with the respective number of edges and minimum degree 3, since adding a degree 2 vertex to a graph always doubles the number of realisations (e.g., [DG24, Lemma 7.1]). The two missing entries have not been computed yet, due to the large number of graphs. For $k=0$ these numbers were computed in [CGG+18].
@@
Table 2: The values of $\mathbf{M}_{2}^{k}(n)$ for $k\in\{0,1,2,3\}$ and $n\leq 11$. The graphs that obtain these numbers can be seen in Table 4.
@@
Similarly, we can get bounds for any $\mathbf{M}_{2}^{k}(n)$ by gluing on a globally rigid subgraph with 2|V|3+k2|V|-3+k edges. Both $\mathbf{M}_{2}^{2}(n)$ and $\mathbf{M}_{2}^{3}(n)$ grow at least as $\left(\sqrt[6]{172}\right)^{n}$, which is approximately $2.35824^{n}$. For $k=3$ this is obtained from the respective graph in Table 4 with 11 vertices. For $k=2$ the graph from the table does not have a globally rigid subgraph with 2|V|3+k2|V|-3+k edges, so we instead use another graph with $c_{2}(G)=172$ (see Appendix B).
@@ 5.2 Computational results for spherical realisation numbers
We have also computed all spherical realisation numbers for rigid graphs with at most 10 vertices ( see [DGS+25]). Additionally for $n\geq 11$ we computed graphs with minimum degree 3 which suffices for attaining the maximum. Table 3 shows the maximal spherical realisation number $\mathbf{M}_{2}^{\circ,k}(n)$ for graphs with $n$ vertices and 2n3+k2n-3+k edges. The realisation numbers for $k=0$ have been previously computed in [GGS20, Gra25].
@@
Table 3: The maximal number of spherical realisations for a given number of vertices $\mathbf{M}_{2}^{\circ,k}(n)$ for $k\in\{0,1,2,3\}$ and $n\leq 10$. Graph that obtain these numbers can be seen in Table 5.
All of the graphs from Table 5 have a suitable globally rigid subgraph on which to use a generalised fan construction. This yields that $\mathbf{M}_{2}^{\circ,1}(n)$ grows at least as $(\sqrt[7]{896})^{n}$, which is approximately $2.64094^{n}$. Both $\mathbf{M}_{2}^{\circ,2}(n)$ and $\mathbf{M}_{2}^{\circ,3}(n)$ grow at least as $\left(\sqrt[5]{128}\right)^{n}$, which is approximately $2.63902^{n}$. These bounds are obtained by the values in Table 3; in particular, the corresponding graphs with 10 vertices in Table 5. In this case the graphs with 11 vertices do not yield a better bound. Again, further computations can improve this bound.
Similarly to [DG24, Section 7], we compare the spherical realisation count to the planar count, but for all rigid graphs. Figure 3 shows this comparison for graphs with 5–9 vertices. That a significant proportion of all graphs have a realisation number of 1 is a consequence of the sharp threshold characterisation for Erdős-Renyi random graph global rigidity [JSS07, LNPR23]. The full data on rigid realisation numbers for up to 10 vertices is available at [DGS+25].

### v2

@@ 1.2 Statement of result
Given Theorem˜1, we can calculate the realisation number for any rigid graph. Indeed, let $G$ be any rigid graph. If $G$ is minimally rigid, then by the existing algorithm of [CGG+18], its realisation number can be computed. If $G$ is globally rigid, its realisation number is $1$. If neither of these occur, we apply Theorem˜1; we can write $c_{2}(G)$ in terms of realisation numbers of graphs on strictly fewer vertices. The algorithm then proceeds recursively; if any of the smaller graphs are minimally rigid or globally rigid, we can compute their realisation number. If not, we apply Theorem˜1 again. This process clearly concludes at some point, since we have a finite number of vertices in $G$ and the number of vertices in the graphs considered strictly decreases at each stage.
It is natural to ask how often Theorem˜1 is needed when computing realisation numbers. Table˜1 shows how many graphs on few vertices are minimally/globally/redundantly rigid. What is particularly interesting is the last column of Table˜1, which shows the number of rigid graphs which are neither minimally rigid nor redundantly rigid and which do not have a 2-cut, i. e., those graphs that require Theorem˜1(i) (the more computationally involved of the two cases) to compute their realisation number.
@@ 1.3 Examples
Figure 1: (Left): The rigid graph $G$ given in Example˜2. (Middle): The maximally rigid subgraphs of $G^{\prime}=G-\{1,2\}$. (Right): The minimally rigid subgraph $H$ of $G$.
@@
Let us consider the graph $G$ from Figure˜1 (left). This graph is clearly not minimally rigid since it has one too many edges. It is not redundantly rigid either and hence not globally rigid, since the deletion of the edge $\{1,2\}$ would result in a non-rigid graph $G^{\prime}$.
We split the graph $G^{\prime}$ into its maximal rigid components $G_{1},\ldots,G_{6}$ as shown in Figure˜1 (middle). Let $G_{1}$ be the graph with vertices $4,5,6,7$. It is the only subgraph that is not minimally rigid. We compute $c_{2}(G_{1})=1$ since it is globally rigid. Each other subgraph $G_{i}$ is a single edge, and hence the respective $H_{i}$ are also single edges. This in turn implies $c_{2}(G_{i})=c_{2}(H_{i})=1$ for each $i\neq 1$. It is not so hard to convince oneself that the graph $H_{1}$ obtained from $G_{1}$ by deleting the edge $\{5,7\}$ is minimally rigid with $c_{2}(H_{1})=2$. The algorithm given in [CGG+18] applied to the graph $H=H_{1}\cup\ldots\cup H_{6}\cup\{1,2\}$ gives that $c_{2}(H)=24$. By applying Theorem˜1(i) we get $c_{2}(G)=c_{2}(H)\prod_{i=1}^{m}\frac{c_{2}(G_{i})}{c_{2}(H_{i})}=24\cdot\frac{1}{2}=12$.
@@
Now consider the graph $\widetilde{G}$ from Figure˜2 (left). Again, this graph is neither minimally rigid nor is it redundantly rigid. The algorithm described in [CGG+18] gives that the graph $\widetilde{H}=\widetilde{G}-\{5,7\}$ has a realisation number of 672. By combining Theorem˜1(i) ($e=\{10,11\}$) with the observation that the subgraph generated on the first 7 vertices is the graph $G$ from Example˜2, we have that $c_{2}(\widetilde{G})=c_{2}(\widetilde{H})\cdot\frac{c_{2}(G)}{c_{2}(H)}=672\cdot\frac{12}{24}=336$.
@@
Figure 2: (Left): The rigid graph $\widetilde{G}$ given in Example˜3. (Middle): The maximally rigid subgraphs of $\widetilde{G}-\{10,11\}$. (Right): The minimally rigid subgraph $\widetilde{H}$ of $\widetilde{G}$.
@@ 1.4 Structure of paper and notation
This paper is structured as follows. We cover all the required background material for rigidity theory in Section˜2. In Section˜3, we provide a proof of Theorem˜1. This algorithm is then adapted in Section˜4 to compute spherical realisation numbers and rank-3 PSD matrix completions. We conclude the paper in Section˜5 with a range of computational results.
@@ 2.2 Algebraic geometry and rigidity
We now wish to view our realisation numbers as degrees of dominant maps of varieties. We note that different authors use various definitions of a ‘realisation variety’, using techniques such as pinning vertices (e.g., [JO19, DG24]). In this paper we do not wish to use pinning, as it makes the upcoming decomposition map (Section˜3) difficult to define. Instead, we want to exploit the following useful property of the variety $R_{V}$: by projecting away coordinates we can reach $R_{V^{\prime}}$ for any $V^{\prime}\subseteq V$.
@@
If $G$ is rigid, then the number of connected components of $f_{G}^{-1}(f_{G}(\rho))$ for generic $\rho$ is $2c_{2}(G)$; see, for example, [LMSW25, Lemma 5]. Similarly, if $G$ is rigid then $\deg(p_{G})=c_{2}(G)$. A proof of this fact can be found in Appendix˜A.
3 Proof of Theorem˜1
As stated previously, Theorem˜1(ii) was originally proven by Jackson and Owen [JO19, Theorem 6.6]. We now proceed with Theorem˜1(i). The main idea of the proof is to factorise the graph map $p_{G}$ in a suitable way, using the following map.
@@ 3 Proof of Theorem˜1
The proof of Lemma˜7 extends to higher dimensions, however the decomposition of $H$ into minimally rigid subgraphs sharing at most one vertex is no longer guaranteed. This is important for the dominance of the decomposition map $d_{S}$; if, in three dimensions, two of the subgraphs, say $H_{1}$ and $H_{2}$ of $G$ share a pair of vertices $v_{1},v_{2}$, then after applying $d_{S}$ to any realisation $\rho$ of $G$, we must have that, for $d_{S}(\rho)=(\rho_{1},\rho_{2},\ldots,\rho_{t})$, the equation $d(\rho_{1}(v_{1}),\rho_{1}(v_{2}))=d(\rho_{2}(v_{1}),\rho_{2}(v_{2}))$ is always satisfied. This shows that $d_{S}$ cannot be a dominant map in such a situation. The dominance of $d_{S}$, however, is crucial for the proof of Theorem˜1. Hence, extending our results to higher dimensions is a subject for further research.
@@
The inclusion $I_{G}\subseteq I_{G_{1}}\times\ldots\times I_{G_{m}}\times\mathbb{C}$ is clear. It now suffices to show that both sets have the same dimension: since both $I_{G}$ and $I_{G_{1}}\times\ldots\times I_{G_{m}}\times\mathbb{C}$ are irreducible and closed in the Zariski topology, this immediately implies the sets are equal. Since $G$ is rigid, we have $\dim I_{G}=r(G)=2|V(G)|-3$. By Lemma˜6, we then have
@@
We are now ready to prove the first case of Theorem˜1.
@@
As in the proof of Lemma˜7, the graph map $p_{G}:R_{V(G)}\to I_{G}$ factors into the product of graph maps of subgraphs
@@
and the decomposition map $d_{S}$; this can be seen in the following commutative diagram, where we note that by Lemma˜9 we have $I_{G}=I_{G_{1}}\times\cdots\times I_{G_{m}}\times I_{e}$:
@@
The map $d_{S}$ is also the decomposition map for the decomposition $\{H_{1},\dots,H_{m},e\}$ of $H$, since $R_{V(G)}=R_{V(H)}$ and $R_{V(G_{i})}=R_{V(H_{i})}$ for all $i$. By Lemma˜5, $H$ is minimally rigid, and by Lemma˜7, $d_{S}$ is a generically finite and dominant map with degree $\frac{c_{2}(H)}{c_{2}(H_{1})\cdots c_{2}(H_{m})}$. For the subgraph $e$, note that $R_{e}=I_{e}=\mathbb{C}$, and $p_{e}$ is the identity map. We then have
@@
Proof of Theorem˜1.
Let $G$ be a graph that is rigid but not globally rigid. By [Hen92], $G$ is either not redundantly rigid, or $G$ is not 3-connected. If $G$ is either not redundantly rigid, then there exists an edge $e\in E(G)$ such that GeG-e is not rigid. Case (i) now follows by Lemma˜10. If $G$ is not 3-connected, we see that Case (ii) holds by [JO19, Theorem 6.6]. ∎
@@ 4.1 Computing spherical realisation numbers
Our proof of Theorem˜1(i) also extends to counting realisations on the (complex) sphere $\mathbb{S}^{2}$. Indeed, a graph is minimally rigid in $\mathbb{C}^{2}$ if and only if is also minimally rigid on the sphere (see for instance [SW07]). Furthermore, a graph is globally rigid in $\mathbb{C}^{2}$ if and only if it is globally rigid in $\mathbb{S}^{2}$; the real variant of this statement follows from a result of Connelly and Whiteley [CW10, Theorem 12], which can then be adapted to the complex setting using a combination of results from Gortler and Thurston [GT14, Theorem 1] and the first and second authors [DG24, Theorem 1.2]. Therefore, Jackson and Jordán’s characterisation of globally rigid graphs also holds in the sphere, and we have the same division into two cases.
@@
The graph from Example˜2 has $c^{\circ}_{2}(G)=16$ since $c^{\circ}_{2}(H)=32$. This is the smallest rigid but not minimally rigid graph where $c_{2}(G)\neq c^{\circ}_{2}(G)$.
@@ 5.1 Computational results for realisation numbers
We focus now on rigid graphs with 2|V|3+k2|V|-3+k edges and a high number of realisations. For $k=0$, i.e. the minimally rigid case, this was done before in [GKT20, Gra25]. Let $\mathbf{M}_{2}^{k}(n)$ be the maximal $c_{2}(G)$ over all rigid graphs $G=(V,E)$ with |E|=2|V|3+k|E|=2|V|-3+k. We show in Table˜2 some values of $\mathbf{M}_{2}^{k}(n)$, which can be found by computing the realisation numbers of all rigid graphs with the respective number of vertices and edges. Note that for $n\leq 10$ we have computed realisation numbers of all rigid graphs which satisfy the desired edge count. They can be found at [DGS+25]. For $n\geq 11$ we only computed realisations numbers for all rigid graphs with the respective number of edges and minimum degree 3, since adding a degree 2 vertex to a graph always doubles the number of realisations (e.g., [DG24, Lemma 7.1]). The two missing entries have not been computed yet, due to the large number of graphs. For $k=0$ these numbers were computed in [CGG+18].
@@
Table 2: The values of $\mathbf{M}_{2}^{k}(n)$ for $k\in\{0,1,2,3\}$ and $n\leq 11$. The graphs that obtain these numbers can be seen in Table˜4.
@@
Similarly, we can get bounds for any $\mathbf{M}_{2}^{k}(n)$ by gluing on a globally rigid subgraph with 2|V|3+k2|V|-3+k edges. Both $\mathbf{M}_{2}^{2}(n)$ and $\mathbf{M}_{2}^{3}(n)$ grow at least as $\left(\sqrt[6]{172}\right)^{n}$, which is approximately $2.35824^{n}$. For $k=3$ this is obtained from the respective graph in Table˜4 with 11 vertices. For $k=2$ the graph from the table does not have a globally rigid subgraph with 2|V|3+k2|V|-3+k edges, so we instead use another graph with $c_{2}(G)=172$ (see Appendix˜B).
@@ 5.2 Computational results for spherical realisation numbers
We have also computed all spherical realisation numbers for rigid graphs with at most 10 vertices ( see [DGS+25]). Additionally for $n\geq 11$ we computed graphs with minimum degree 3 which suffices for attaining the maximum. Table˜3 shows the maximal spherical realisation number $\mathbf{M}_{2}^{\circ,k}(n)$ for graphs with $n$ vertices and 2n3+k2n-3+k edges. The realisation numbers for $k=0$ have been previously computed in [GGS20, Gra25].
@@
Table 3: The maximal number of spherical realisations for a given number of vertices $\mathbf{M}_{2}^{\circ,k}(n)$ for $k\in\{0,1,2,3\}$ and $n\leq 10$. Graph that obtain these numbers can be seen in Table˜5.
All of the graphs from Table˜5 have a suitable globally rigid subgraph on which to use a generalised fan construction. This yields that $\mathbf{M}_{2}^{\circ,1}(n)$ grows at least as $(\sqrt[7]{896})^{n}$, which is approximately $2.64094^{n}$. Both $\mathbf{M}_{2}^{\circ,2}(n)$ and $\mathbf{M}_{2}^{\circ,3}(n)$ grow at least as $\left(\sqrt[5]{128}\right)^{n}$, which is approximately $2.63902^{n}$. These bounds are obtained by the values in Table˜3; in particular, the corresponding graphs with 10 vertices in Table˜5. In this case the graphs with 11 vertices do not yield a better bound. Again, further computations can improve this bound.
Similarly to [DG24, Section 7], we compare the spherical realisation count to the planar count, but for all rigid graphs. Figure˜3 shows this comparison for graphs with 5–9 vertices. That a significant proportion of all graphs have a realisation number of 1 is a consequence of the sharp threshold characterisation for Erdős-Renyi random graph global rigidity [JSS07, LNPR23]. The full data on rigid realisation numbers for up to 10 vertices is available at [DGS+25].